OpenStudy (magbak):

PLEASE HELP ME I WILL AWARD MEDAL!!!!!!!!!!!!!!!!! Apply the formula of a rectangular box (V = lwh) to find the volume of the object. Now suppose you knew the volume of this object and the relation of the length to the width and height, but did not know the length. Rewriting the equation with one variable would result in a polynomial equation that you could solve to find the length. Rewrite the formula using the variable x for the length. Substitute the value of the volume found in step 2 for V and express the width and height of the object in terms of x plus or minus a constant

OpenStudy (magbak):

@whpalmer4 @Hero

OpenStudy (magbak):

@jim_thompson5910

OpenStudy (anonymous):

Since your measurements are Height: 4 Length: 11 Width:7 multiply them together to get V = LWH V = 11*7*4 V = 308 So the volume is 308 cubic units

OpenStudy (anonymous):

ok so you know the volume is 308 cubic units now let's say you either forgot the length, or you only told your friends the width and height so the length is unknown and it's x for now V = LWH 308 = x*7*4 308 = 28x 28x = 308 I'll let you finish

OpenStudy (anonymous):

for part 3 you let x = length and you let either the width or height be unknow then you say something like if x = length and the width is unknown, but you know it's 4 units less than the length, then you can say length = x width = x-4 height = 4 and you can plug all this info into V = LWH V = LWH 308 = x(x-4)*4 308 = 4x^2 - 16x 4x^2 - 16x - 308 = 0 and you would solve this for x to get the length, which you would later use to find the width

OpenStudy (magbak):

Thank you do very much.

OpenStudy (anonymous):

No problem :)

OpenStudy (magbak):

@heather040200

OpenStudy (anonymous):

yo

OpenStudy (magbak):

I sent you a message.

OpenStudy (magbak):

Apply the formula of a rectangular box (V = lwh) to find the volume of the object. Now suppose you knew the volume of this object and the relation of the length to the width and height, but did not know the length. Rewriting the equation with one variable would result in a polynomial equation that you could solve to find the length. Rewrite the formula using the variable x for the length. Substitute the value of the volume found in step 2 for V and express the width and height of the object in terms of x plus or minus a constant Dimensions of 7 inches x 7 inches x 6 inches

OpenStudy (anonymous):

I believe you use the same formula V=LWH

OpenStudy (magbak):

No problem

OpenStudy (magbak):

Where did you go now. @heather040200

OpenStudy (anonymous):

yo

OpenStudy (magbak):

you left again with out helping me.

OpenStudy (anonymous):

I said I believe you use the same formula V=LWH just plg in the numbers

OpenStudy (magbak):

No I need the last part with the standard form I know that volume it is 294 I just need the last equation PLEASE :}

OpenStudy (anonymous):

oooh ok I was confused LOL

OpenStudy (magbak):

{:[]}